A projectile is projected with speed u at an angle of 60 degrees with the horizontal from the foot of an inclined plane.If the projectile hits the plane horizontally,the range on the inclined plane will be:


A)((u^2)*underroot 21)/2g


B)3u^2/4g


C)u^2/2g


D)underroot 21*u^2/8g


ANS:D

Grade:12th Pass

1 Answers

Akash Kumar Dutta
11 years ago

Dear Sameer,
time = 1/2 . time of flight=usin(o)/g=u.sin60/g
x axis distance = ucos60.t=ucos60.usin60/g=(u^2)sin60.cos60/g
distance on y=H(max)=(u^2)sin^2 (60)/2g
Hence total displacement=Range on inclined plane=root(x^2+y^2)=root(21)u^2/8g
(ANS)

Regards.